# compat=1.1.0

### 19 Dec 2023 v1

@@ Abstract
In this paper, we present predictions for scattering cross-section the of Higgs boson pair production via photon fusion at future muon colliders, focusing specifically the production processes $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$. We investigated the impact of three choices the photon structure functions on cross-section predictions for a range model input parameters within the theoretical framework of the Higgs Triplet Model[1, 2, 3, 4].
@@ I Introduction
Double Higgs production is important for testing the Higgs self-coupling [5] which is responsible for providing mass to elementary particles and the shape of the Higgs potential. In the SM, the Higgs potential for the Higgs field $\phi$ is defined as,
$V(\phi)=\mu^{2}\phi\phi^{*}+\frac{\lambda^{2}}{2}(\phi^{*}\phi)^{2}\,$ (1)
where $\lambda>0$ and $\mu^{2}<0$. The minimum value of the Higgs potential occurs at $v^{2}=-\mu^{2}/\lambda$. Following spontaneous symmetry breaking, the Higgs boson acquires a mass $M_{H^{0}}=-2\mu^{2}=2\lambda v^{2}$. In the Standard Model, the relationships among the physical Higgs mass, cubic interaction ($\lambda_{hhh}$), and quartic interaction ($\lambda_{hhhh}$) are uniquely defined and can be expressed as $\lambda_{hhh}=v\lambda_{hhhh}=3M_{H}^{2}/v$. However, the trilinear Higgs coupling is challenging to measure directly, as it requires the production of two or more Higgs bosons simultaneously [6, 7, 8]. Also, to measure the trilinear Higgs coupling at the LHC requires high luminosity because processes that involve the trilinear Higgs coupling are rare in SM [7]. The anticipated cross section for the production of Higgs boson pairs via gluon gluon fusion stands at approximately $36.69\ \mathrm{fb}$ at a center of mass energy of 14 TeV [9, 10]. Even with the highest possible enhancement of both center of mass energy and integrated luminosity at LHC, the accurate extraction of $\lambda_{hhhh}$ remains a formidable challenge. Nevertheless, there is optimism regarding the feasible observation of Higgs boson pair production and the determination of $\lambda_{hhh}$. However, achieving these goals might necessitate an integrated luminosity of $3000\ \mathrm{fb}^{-1}$ at the HL-LHC, [9, 11, 12]. The production of Higgs boson pairs via gluon fusion has been studied across various theoretical frameworks, including the Standard Model [13], Minimal Supersymmetric Standard Model [14], Lee Wick Standard Model [15], and Singlet Extension Model [16].
Muon colliders offer several advantages over proton colliders, which can help to avoid some of the challenges associated with measuring the trilinear Higgs coupling. Muon colliders can reach higher center of mass energies than proton colliders ($\sqrt{s}\sim\mathcal{O}(10\ \mathrm{TeV})$), which can increase the production rate of triple Higgs boson events[17]. Here we want to analyse the double Higgs production initiated by collinear photons radiated by high energy muon beams. The spectrum of photons with an energy fraction $x$ emitted by a charged lepton with an initial energy $E$ is described by the Weizsäcker-Williams spectrum, also known as the Leading Order Effective Photon Approximation (EPA) [18, 19],
@@
The splitting functions are, $P_{\gamma,l}=(1+(1-x^{2})/x$ for $l\rightarrow{\gamma}$, and $P_{l/l}(x)=(1+x^{2})/(1-x)$ for $l\rightarrow{l}$.
Double Higgs boson production via photon fusion has been examined in the SM at multi-TeV muon collider [17] and in the Two Higgs Doublet Model (2HDM) at the International Linear Collider (ILC) [20]. However, these models do not account for the impact of the $H^{\pm\pm}$ boson’s involvement in the $\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$ processes. This aspect is addressed within the context of the Higgs Triplet Model (HTM). In the HTM, a noteworthy hierarchy is observed between the masses of $H^{\pm\pm}$ and $H^{\pm}$, and this influence is thoroughly investigated in the context of this paper.
Initially, we examine the partonic cross sections involving $\gamma\gamma$ fusion leading to the production of Higgs pairs in the HTM. Following that, we numerically compute the cross-sections for $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$ by performing the convolution of the parton distribution functions (PDFs) with the partonic scattering cross-section. We consider the muon collider at a benchmark energy of $\sqrt{s}=3$ TeV, with the integrated luminosity [21] scaled according to
@@
This paper is organized as follows: In Section II, we present a description of the Higgs Triplet model, while Section III provides constraints on the model parameters and input values. Section IV outlines the computational method used for calculating the scattering amplitude and cross sections. Moving to Section V, we conduct an analysis and present numerical results encompassing partonic cross sections, collider simulations for $h^{0}h^{0}$ and $A^{0}A^{0}$ production, along with the discussion. Additionally, we explore the examined couplings, particularly addressing the decoupling and weak-coupling limits that may arise in the calculations. We summarize our findings in Section VI.
@@ II Theoretical Framework
Our theoretical framework is built upon the HTM, also known as the Type II Seesaw [1, 2]. In the HTM, alongside the Standard Model weak doublet, represented as $\Phi\sim(1,2,1/2)$, there is an additional Higgs triplet denoted as $\Delta\sim(1,3,2)$ which transforms under the $SU(2)_{L}$ gauge group. The motivation for the Type II Seesaw Model stems from the observation that two doublets can be decomposed into a triplet and a singlet representation ($2\otimes 2=3\oplus 1$). It is assumed that these additional fields possess a mass scale that is significantly higher than the electroweak scale. By introducing these extra scalar fields, new Yukawa couplings are established between the SM lepton fields and the scalar fields. These Yukawa couplings generate relatively small Majorana neutrino masses without requiring right handed neutrinos. Due to the higher mass scale of the additional scalar fields, the resulting masses for the neutrinos are much smaller compared to the electroweak scale. In addition to the Yukawa interactions present in the Standard Model, the Yukawa sector of the Type II Seesaw Model incorporates interactions between the Higgs triplet field ($\Delta$) and the lepton fields:
@@
The Higgs Triplet Model incorporates five dimensionless parameters ($\lambda$, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$), two real parameters with mass dimensions ($\mu_{\Phi}$, $\mu_{\Delta}$), and a lepton number violating parameter with positive mass dimension ($\mu$). The vacuum expectation value ($v_{\Phi}$) of the Higgs field, which is responsible for breaking the electroweak symmetry, takes place in a direction that does not introduce any electric charge or alter the electrically neutral nature of the vacuum. The Higgs triplet field acquires a non-zero vacuum expectation value ($v_{\Delta}$) that leads to the spontaneous breaking of the electroweak symmetry. Consequently, the neutral components of the Higgs triplet field develop non-zero vacuum expectation value, while the charged components remain zero.
In the absence of CP violation, the scalar fields $\phi^{0}$ and $\delta^{0}$ can be parameterized as follows: $\phi^{0}=\frac{1}{\sqrt{2}}(v_{\Phi}+\phi+i\chi)$ and $\delta^{0}=\frac{1}{\sqrt{2}}(v_{\Delta}+\delta+i\eta)$. This parameterization involves shifting the real part of $\phi^{0}$, denoted as $\phi$, and the real part of $\delta^{0}$, denoted as $\delta$, around their vacuum expectation values (VEVs). As a result, a $10\times 10$ squared mass matrix is obtained to describe the scalars in the model. After diagonalizing the mass matrix and utilizing the the following rotation matrices:
@@
By expressing the couplings ($\lambda$,$\lambda_{1,2,3,4}$ and $\mu$) in terms of physical Higgs masses, the mixing angle $\alpha$ and VEVs ($v_{\Delta}$ and $v_{\Phi}$), we can directly relate the strength of the interactions to the masses of the particles involved and the vacuum expectation values that characterize the symmetry breaking. Furthermore, through this consistent parameterization of the couplings, we can readily compare the predictions and implications of the HTM with those of other models and experimental observations. Now, using Eq. (9), (10) and (11), one can obtain the following relations:
@@
By utilizing these parameter spaces, computations can be performed in a bidirectional manner, allowing for the evaluation of the quantities and relations within the HTM using either set of parameters. Moreover, $v_{\Delta}$ and $v_{\Phi}$ are re-parameterized following the conventions of Ref. [1]:
@@ III Constraints on the potential
In the HTM, vacuum stability is required to ensure that the electroweak symmetry breaking minimum of the Higgs potential is stable and that the vacuum dies decay into a lower-energy state. In order to ensure that the scalar potential $V(\Phi,\Delta)$ in the HTM is always bounded from below (BFB) and does not lead to instability of the vacuum state, we must impose the following constraints on the model parameters from Ref. [1]:
@@ IV The Details of the Calculation
To calculate the total cross sections, we employed the Weizsäcker-Williams approximation, which represents the Leading Order (LO) contribution for the photon PDFs, as shown in Eq. (2). In the work presented in Ref. [23], the solution to the DGLAP equations has been achieved through iterative techniques. As a second approach in performing the cross section calculations, we use their second order corrected PDF, denoted as LO$+\mathcal{O}(\alpha_{e}^{2}t^{2})$ as well, see the Eq. (53). Thirdly, we used EW PDF at the Leading -Log (LL) accuracy available at [24].
In this paper, the explicit presentation of the matrix element expressions has been omitted due to their length. We implemented the HTM Lagrangian into FeynRules [^1] Even though there is a FeynRules model file available at Ref. [33], we have implemented our own FeynRules model file.. The generation of the FeynArts models files was performed by FeynRules [26]. The generation of one-loop amplitudes was performed by FeynArts [27] and the subsequent generation of the matrix element squared was performed using FormCalc [28]. In order to incorporate photon structure functions, distinct Fortran subroutines were developed for both LO and the second order corrected EPA. The numerical assessments of the integration over the $2\rightarrow 2$ phase space were carried out using the CUBA library[29, 30].
@@ V Analysis, Results and Discussion
Initially, we examine the overall rates of the partonic processes $\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$ in the center of mass system of $\gamma\gamma$ before convoluting them with the photon energy spectrum in a muon collider. In our numerical analysis, we employed $M_{W}=80.45\ \mathrm{GeV}$, $M_{Z}=91.18\ \mathrm{GeV}$, the Weinberg angle $S_{W}=\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}$, and the fine structure constant $\alpha_{e}=1/137.03598$.
@@
    Utilizing input values of $M_{H^{0}}=463.5$ GeV, $M_{H^{\pm\pm}}=428.31$ GeV, $v_{\Phi}=245.9$ GeV, $v_{\Delta}=1$ GeV, and $\cos{\alpha}=0.999$, we determined the maximum value of $M_{A^{0}}$ to be approximately 463.5 GeV, with a corresponding maximum value of $M_{H^{\pm}}$ at 446.28 GeV. This parameter set was employed to compute partonic cross-sections for the scenario where $\Delta M>0$.
@@ V.1 $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$ Cross sections
Figure 12: Relative difference of the Higgs boson pair production cross section with HTM trilinear coupling and the SM cross section as a function of trilinear $H^{0}h^{0}h^{0}$ coupling $\sqrt{s}=3\ \mathrm{TeV}$.
Figure 13: Relative difference of the Higgs boson pair production cross section with HTM trilinear coupling and the SM cross section as a function of trilinear $H^{0}h^{0}h^{0}$ coupling $\sqrt{s}=3\ \mathrm{TeV}$.
Figure 14: Relative difference of the Higgs boson pair production cross section with HTM trilinear coupling and the SM cross section as a function of trilinear $H^{0}h^{0}h^{0}$ coupling $\sqrt{s}=3\ \mathrm{TeV}$.
Table 1: The presented mass ranges in this table correspond to the input values $\tan{\beta}=0.001$, $\lambda=0.51$, $\lambda_{1}=10$, $\lambda_{2}=1$, $\lambda_{3}=-1$, and $\mu=[0.2,1.5]$. The numerical values demonstrate that $\Delta M>0$ for $\lambda_{4}>0$, $\Delta M=0$ for $\lambda_{4}=0$, and $\Delta M<0$ for $\lambda_{4}<0$.
@@
    According to Figs. 6, 7 and 8, the cross section for $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0}$ at $M_{H^{\pm}}\approx M_{H^{\pm\pm}}$ is approximately $10$ times larger than the cross section when $M_{H^{\pm\pm}}>M_{H^{\pm}}$, and it is approximately $60$ times larger when $M_{H^{\pm}}>M_{H^{\pm\pm}}$ compared to $M_{H^{\pm\pm}}\approx M_{H^{\pm}}$.
@@
    In Figs. 910 and 11, the cross section for $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow A^{0}A^{0}$ is notably higher, being approximately $1.2$ times larger at $M_{H^{\pm}}>M_{H^{\pm\pm}}$ compared to the scenario where $M_{H^{\pm}}\approx M_{H^{\pm\pm}}$. Similarly, when $M_{H^{\pm}}>M_{H^{\pm\pm}}$, the cross section is approximately $5.6$ times larger than the case when $M_{H^{\pm}}<M_{H^{\pm\pm}}$.

### 05 Apr 2024 v2

@@ Abstract
In this paper, we present predictions for scattering cross section the of Higgs boson pair production via photon fusion at future muon colliders, focusing specifically the production processes $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$. We investigated the impact of three choices the photon structure functions on cross-section predictions for a range model input parameters within the theoretical framework of the Higgs triplet model[1, 2, 3, 4].
@@ I Introduction
Double Higgs production is important for testing the Higgs self-coupling [5, 6] which is responsible for providing mass to elementary particles and the shape of the Higgs potential. In the SM, the Higgs potential for the Higgs field $\phi$ is defined as,
$V(\phi)=\mu^{2}\phi\phi^{*}+\frac{\lambda}{2}(\phi^{*}\phi)^{2}\,$ (1)
where $\lambda>0$ and $\mu^{2}<0$. The minimum value of the Higgs potential occurs at $v^{2}=-\mu^{2}/\lambda$. Following spontaneous symmetry breaking, the Higgs boson acquires a mass $M_{H}=-2\mu^{2}=2\lambda v^{2}$. In the Standard Model, the relationships among the physical Higgs mass, cubic interaction ($\lambda_{hhh}$), and quartic interaction ($\lambda_{hhhh}$) are uniquely defined and can be expressed as $\lambda_{hhh}=v\lambda_{hhhh}=3M_{H}^{2}/v$. However, the trilinear Higgs coupling is challenging to measure directly, as it requires the production of two or more Higgs bosons simultaneously [7, 8, 9]. Also, to measure the trilinear Higgs coupling at the LHC requires high luminosity because processes that involve the trilinear Higgs coupling are rare in SM [8]. The anticipated cross section for the production of Higgs boson pairs via gluon gluon fusion stands at approximately $36.69\ \mathrm{fb}$ at a center of mass energy of 14 TeV [10, 11]. Even with the highest possible enhancement of both center of mass energy and integrated luminosity at LHC, the accurate extraction of $\lambda_{hhhh}$ remains a formidable challenge. Nevertheless, there is optimism regarding the feasible observation of Higgs boson pair production and the determination of $\lambda_{hhh}$. However, achieving these goals might necessitate an integrated luminosity of $3000\ \mathrm{fb}^{-1}$ at the HL-LHC, [10, 12, 13]. The production of Higgs boson pairs via gluon fusion has been studied across various theoretical frameworks, including the Standard Model [14], minimal supersymmetric Standard Model [15, 16], Lee Wick Standard Model [17], and singlet extension model [18].
Muon colliders offer several advantages over proton colliders, which can help to avoid some of the challenges associated with measuring the trilinear Higgs coupling. Muon colliders can reach higher center of mass energies than proton colliders [$\sqrt{s}\sim\mathcal{O}(10\ \mathrm{TeV})$], which can increase the production rate of triple Higgs boson events[19]. Here we want to analyze the double Higgs production initiated by collinear photons radiated by high energy muon beams. The spectrum of photons with an energy fraction $x$ emitted by a charged lepton with an initial energy $E$ is described by the Weizsäcker-Williams spectrum, also known as the leading order effective photon approximation (EPA) [20, 21],
@@
The splitting functions are, $P_{\gamma,l}=(1+(1-x^{2}))/x$ for $l\rightarrow{\gamma}$, and $P_{l/l}(x)=(1+x^{2})/(1-x)$ for $l\rightarrow{l}$.
Double Higgs boson production via photon fusion has been examined in the SM at multi-TeV muon collider [19] and in the Two Higgs doublet model (2HDM) at the International Linear Collider (ILC) [22, 23, 24]. However, these models do not account for the impact of the $H^{\pm\pm}$ boson’s involvement in the $\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$ processes. This aspect is addressed within the context of the Higgs triplet model (HTM). In the HTM, a noteworthy hierarchy is observed between the masses of $H^{\pm\pm}$ and $H^{\pm}$, and this influence is thoroughly investigated in the context of this paper.
Initially, we examine the partonic cross sections involving $\gamma\gamma$ fusion leading to the production of Higgs pairs in the HTM. Following that, we numerically compute the cross sections for $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$ by performing the convolution of the parton distribution functions (PDFs) with the partonic scattering cross section. We consider the muon collider at a benchmark energy of $\sqrt{s}=3$ TeV, with the integrated luminosity [25] scaled according to
@@
This paper is organized as follows: In Sec. II, we present a description of the Higgs triplet model, while Sec. III provides constraints on the model parameters and input values. Sec. IV outlines the computational method used for calculating the scattering amplitude and cross sections. Moving to Sec. V, we conduct an analysis and present numerical results encompassing partonic cross sections, collider simulations for $h^{0}h^{0}$ and $A^{0}A^{0}$ production, along with the discussion. Additionally, we explore the examined couplings, particularly addressing the decoupling and weak-coupling limits that may arise in the calculations. We summarize our findings in Sec. VI.
@@ II Theoretical Framework
Our theoretical framework is built upon the HTM, also known as the type II seesaw [1, 2, 3, 4]. In the HTM, alongside the Standard Model weak doublet, represented as $\Phi\sim(1,2,1/2)$, there is an additional Higgs triplet denoted as $\Delta\sim(1,3,2)$ which transforms under the $SU(2)_{L}$ gauge group. The motivation for the type II seesaw Model stems from the observation that two doublets can be decomposed into a triplet and a singlet representation ($2\otimes 2=3\oplus 1$). It is assumed that these additional fields possess a mass scale that is significantly higher than the electroweak scale. By introducing these extra scalar fields, new Yukawa couplings are established between the SM lepton fields and the scalar fields. These Yukawa couplings generate relatively small Majorana neutrino masses without requiring right handed neutrinos. Due to the higher mass scale of the additional scalar fields, the resulting masses for the neutrinos are much smaller compared to the electroweak scale. In addition to the Yukawa interactions present in the Standard Model, the Yukawa sector of the type II seesaw model incorporates interactions between the Higgs triplet field ($\Delta$) and the lepton fields:
@@
The Higgs triplet model incorporates five dimensionless parameters ($\lambda$, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$), two real parameters with mass dimensions ($\mu_{\Phi}$, $\mu_{\Delta}$), and a lepton number violating parameter with positive mass dimension ($\mu$). The vacuum expectation value ($v_{\Phi}$) of the Higgs field, which is responsible for breaking the electroweak symmetry, takes place in a direction that does not introduce any electric charge or alter the electrically neutral nature of the vacuum. The Higgs triplet field acquires a nonzero vacuum expectation value ($v_{\Delta}$) that leads to the spontaneous breaking of the electroweak symmetry. Consequently, the neutral components of the Higgs triplet field develop nonzero vacuum expectation value, while the charged components remain zero.
In the absence of CP violation, the scalar fields $\phi^{0}$ and $\delta^{0}$ can be parametrized as follows: $\phi^{0}=\frac{1}{\sqrt{2}}(v_{\Phi}+\phi+i\chi)$ and $\delta^{0}=\frac{1}{\sqrt{2}}(v_{\Delta}+\delta+i\eta)$. This parameterization involves shifting the real part of $\phi^{0}$, denoted as $\phi$, and the real part of $\delta^{0}$, denoted as $\delta$, around their vacuum expectation values (VEVs). As a result, a $10\times 10$ squared mass matrix is obtained to describe the scalars in the model. After diagonalizing the mass matrix and utilizing the following rotation matrices:
@@
By expressing the couplings ($\lambda$,$\lambda_{1,2,3,4}$ and $\mu$) in terms of physical Higgs masses, the mixing angle $\alpha$ and VEVs ($v_{\Delta}$ and $v_{\Phi}$), we can directly relate the strength of the interactions to the masses of the particles involved and the vacuum expectation values that characterize the symmetry breaking. Furthermore, through this consistent parametrization of the couplings, we can readily compare the predictions and implications of the HTM with those of other models and experimental observations. Now, using Eq. (9), (10) and (11), one can obtain the following relations:
@@
By utilizing these parameter spaces, computations can be performed in a bidirectional manner, allowing for the evaluation of the quantities and relations within the HTM using either set of parameters. Moreover, $v_{\Delta}$ and $v_{\Phi}$ are reparametrized following the conventions of Ref. [1]:
@@ III Constraints on the potential
In the HTM, vacuum stability is required to ensure that the electroweak symmetry breaking minimum of the Higgs potential is stable and that the vacuum dies decay into a lower-energy state. In order to ensure that the scalar potential $V(\Phi,\Delta)$ in the HTM is always bounded from below (BFB) and does not lead to instability of the vacuum state, we must impose the following constraints on the model parameters from Ref. [1, 26]:
@@ IV The Details of the Calculation
To calculate the total cross sections, we employed the Weizsäcker-Williams approximation, which represents the leading order (LO) contribution for the photon PDFs, as shown in Eq. (2). In the work presented in Ref. [28], the solution to the DGLAP equations has been achieved through iterative techniques. As a second approach in performing the cross section calculations, we use their second order corrected PDF, denoted as LO$+\mathcal{O}(\alpha_{e}^{2}t^{2})$ as well, see the Eq. (87). Third, we used EW PDF at the leading -log (LL) accuracy available at [29].
In this paper, the explicit presentation of the matrix element expressions has been omitted due to their length. We implemented the HTM Lagrangian into FeynRules [^1] Even though there is a FeynRules model file available at Ref. [40], we have implemented our own FeynRules model file.. The generation of the FeynArts models files was performed by FeynRules [31, 32]. The generation of one-loop amplitudes was performed by FeynArts [33] and the subsequent generation of the matrix element squared was performed using FormCalc [34, 35]. In order to incorporate photon structure functions, distinct Fortran subroutines were developed for both LO and the second order corrected EPA. The numerical assessments of the integration over the $2\rightarrow 2$ phase space were carried out using the CUBA library[36, 37].
@@ V Analysis, Results and Discussion
Initially, we examine the overall rates of the partonic processes $\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$ in the center of mass system of $\gamma\gamma$ before convoluting them with the photon energy spectrum in a muon collider. In our numerical analysis, we employed $M_{W}=80.379\ \mathrm{GeV}$, $M_{Z}=91.18\ \mathrm{GeV}$, the Weinberg angle $S_{W}=\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}$, and the fine structure constant $\alpha_{e}=1/137.03598$.
@@
    Utilizing input values of $M_{H^{0}}=463.5$ GeV, $M_{H^{\pm\pm}}=428.31$ GeV, $v_{\Phi}=245.9$ GeV, $v_{\Delta}=1$ GeV, and $\cos{\alpha}=0.999$, we determined the maximum value of $M_{A^{0}}$ to be approximately 463.5 GeV, with a corresponding maximum value of $M_{H^{\pm}}$ at 446.28 GeV. This parameter set was employed to compute partonic cross sections for the scenario where $\Delta M>0$.
@@ V.1 $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$ cross sections
Figure 12: The relative difference between the cross sections of Higgs boson pair production in the HTM and the SM, as a function of the trilinear scalar coupling $H^{0}h^{0}h^{0}$, is analyzed at a center-of-mass energy of $\sqrt{s}=3\ \mathrm{TeV}$.
Figure 13: The relative difference between the cross sections of Higgs boson pair production in the HTM and the SM, as a function of the trilinear scalar coupling $H^{0}h^{0}h^{0}$, is analyzed at a center-of-mass energy of $\sqrt{s}=3\ \mathrm{TeV}$.
Figure 14: The relative difference between the cross sections of Higgs boson pair production in the HTM and the SM, as a function of the trilinear scalar coupling $H^{0}h^{0}h^{0}$, is analyzed at a center-of-mass energy of $\sqrt{s}=3\ \mathrm{TeV}$.
Table 1: The presented mass ranges in this table correspond to the input values $\tan{\beta}=0.001$, $\lambda=0.51$, $\lambda_{1}=10$, $\lambda_{2}=1$, $\lambda_{3}=-1$, and $\mu=[0.2,2]$. The numerical values demonstrate that $\Delta M>0$ for $\lambda_{4}>0$, $\Delta M=0$ for $\lambda_{4}=0$, and $\Delta M<0$ for $\lambda_{4}<0$.
@@
    According to Figs. 6, 7, and 8, the cross section for $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0}$ at $M_{H^{\pm}}\approx M_{H^{\pm\pm}}$ is approximately $10$ times larger than the cross section when $M_{H^{\pm\pm}}>M_{H^{\pm}}$, and it is approximately $60$ times larger when $M_{H^{\pm}}>M_{H^{\pm\pm}}$ compared to $M_{H^{\pm\pm}}\approx M_{H^{\pm}}$.
@@
    In Figs. 910, and 11, the cross section for $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow A^{0}A^{0}$ is notably higher, being approximately $1.2$ times larger at $M_{H^{\pm}}>M_{H^{\pm\pm}}$ compared to the scenario where $M_{H^{\pm}}\approx M_{H^{\pm\pm}}$. Similarly, when $M_{H^{\pm}}>M_{H^{\pm\pm}}$, the cross section is approximately $5.6$ times larger than the case when $M_{H^{\pm}}<M_{H^{\pm\pm}}$.